# Decoding the AI Pen: Techniques and Challenges in Detecting AI-Generated Text

### v2

@@ Abstract.
LLM, AI-generated Text Detection; Responsible AI; Watermarking; Paraphrasing Attacks; Data Poisoning
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671463isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Natural language generationccs: Computing methodologies Neural networksccs: Security and privacy Social engineering attacks
@@ 1. Introduction
Large Language Models (LLMs) constitute a transformative advancement in the field of natural language processing (NLP). Their applications traverse a wide spectrum of domains, including question answering (Zaib et al., 2021; Bhat et al., 2023), sentiment analysis (Batra et al., 2021; Kheiri and Karimi, 2023) and specially text generation (Senadeera and Ive, 2022). As LLMs are trained on extensive textual corpora, they demonstrate a remarkable capacity to produce human-like text.
@@
However, the ubiquity of LLMs brings forth a convergence of opportunities and challenges that necessitate prudent examination. Among these challenges, we encounter the potential for LLMs to produce biased, toxic, or harmful content (Kuchnik et al., 2023; Deshpande et al., 2023; Weidinger et al., 2021; Wen et al., 2023). Additionally, there are concerns related to intellectual property rights infringement (Peng et al., 2023; Stokel-Walker, 2022) and the misuse of LLMs for malicious purposes, such as disseminating misleading information and propaganda (Vykopal et al., 2023; Mozes et al., 2023). These multifaceted considerations highlight the need for judicious evaluation and ethical handling of LLMs in various contexts.
Leveraging AI-generated text could be regarded as an efficacious strategy for alleviating the aforementioned challenges. However, the convergence of AI-generated content with human-written texts has reached a point where discerning between the two has become increasingly intricate. The task of distinguishing LLM-generated text from human-written content presents a dual challenge. On one hand, identifying disparities can enhance the quality of AI-generated material. On the other hand, this endeavor also complicates the identification process. In recent years, researchers have proposed various methodologies for detecting AI-generated text . These efforts contribute to the ongoing exploration of this landscape.
@@
The rest of this paper is organized as follows: In section 2, we discuss the potential risks arising from the misuse of LLMs. Section 3 provides a comprehensive categorization of various text detection strategies, emphasizing their role as mitigation techniques. Subsequently, in Section 4, we describe the vulnerabilities inherent in these strategies. Moving forward, Section 4.3 engages in a theoretical exploration of the feasibility of detection, followed by section 5 where we propose new avenues of research. Finally, in section 6, we conclude. Figure 1 demonstrates an overview of the paper.
@@ 2.1. Discrimination, Toxicity, and Harms
A recent study by Wen et al. (Wen et al., 2023) delves into the phenomenon of LLMs producing implicit toxic content that remains elusive to conventional toxicity classifiers. The proposed approach leverages reinforcement learning to uncover and expose this implicit toxicity, emphasizing the need to fine-tune classifiers using annotated examples derived from the attack method for enhanced detection capabilities.
@@ 2.2. Factual Inconsistency and Unreliability of AI Responses
To mitigate such mistakes, various strategies have been proposed through fine-tuning (Lewkowycz et al., 2022; Rajani et al., 2019; Zelikman et al., 2022), prompt engineering techniques such as verification, scratchpads  (Cobbe et al., 2021; Nye et al., 2022), Chain-of-Thought (CoT) (Wei et al., 2022), Reversing Chain-of-Thought (RCoT)  (Xue et al., 2023), RLHF (Ziegler et al., 2019; Christiano et al., 2017), iterative self-reflection (Shinn et al., 2023; Madaan et al., 2023), self-consistency (Wang et al., 2023c), society of minds strategy (Du et al., 2023), pruning truthful datasets (Christiano et al., 2023), adjusting the system parameters to limit model creativity (Muneeswaran et al., 2023), external knowledge retrieval (Guu et al., 2020) and training-free methods with likelihood estimation (Kadavath et al., 2022).
Indeed, the existing body of research emphasize the unspeakable capabilities of LLMs, while simultaneously acknowledging their susceptibility to errors. Therefore, it is crucial to approach LLM-generated outputs with caution.
@@ 4. Vulnerabilities of Detection Techniques
Despite various methods mentioned above, most categories of detection strategies are susceptible to paraphrasing or spoofing attacks. To tackle this, retrieval-based detectors have been employed as detection strategies but these detectors store the output of LLMs in a database and perform semantic searches to extract the best matches. Although it enhances the detector’s resilience against paraphrasing attacks, there are privacy concerns related to storing user-LLM conversations. Additionally, this technique is ineffective when dealing with recursive paraphrasing (Sadasivan et al., 2024). Moreover, researchers have discovered that by effectively optimizing prompts, LLMs can successfully evade many detection techniques.
For example, in a recent work (Lu et al., 2023), Lu et al. propose a novel method called Substitution-based In-Context example Optimization (SICO). In SICO, discriminating features are extracted from both human and AI-generated texts. These features, along with a paraphrasing prompt, are combined and fed to the LLM to modify its output. The prompt is carefully optimized through word and sentence replacements, aiming to minimize the chances of detection while maximizing the similarity between human and AI-generated texts. Evaluation results strongly demonstrate the vulnerability of existing methods.
@@ 4.2. Possibility via Sample Abundance
Interestingly, Chakraborty et al. demonstrate that the AUROC curve, as proposed by Sadasivan et al., might be overly conservative for practical detection scenarios. Specifically, they introduce a hidden possibility by replacing $TV(\mathcal{M},\mathcal{H})$ with $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ in AUROC equation, where $m^{\bigotimes n}:=m\bigotimes m\bigotimes\dots\bigotimes m$ (n times) denotes the product distribution over sample set $\mathcal{S}:=\{s_{i}\},i\in\{1,\dots n\}$, as does $h^{\bigotimes n}$. For any detector $\mathcal{D}$, with a given collection of i.i.d. samples either from human or AI, the AUROC curve is:
@@
where $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})=1-exp(-nl_{c}(m,h)+o(n)$ and $l_{c}(m,h)$ is the Chernoff information. Since $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ is an increasing sequence, it eventually converges to $1$ as the number of samples for each distribution increases. Therefore, the upper bound of AUROC increases exponentially with respect to the number of samples. It is clear that if this happens, the total variation distance approaches $1$ quickly, and hence increasing the AUROC. Interested readers are encouraged to refer to (Chakraborty et al., 2023) for further details.
@@ 4.3. Impossibility of Strong Watermarking
5. New Opportunities and Future Research
@@
  • Curating diverse and representative datasets is crucial for training and evaluating detection models. The existing datasets may not comprise all the potential types and sources of AI text. Specially, it is essential to curate a dataset containing outputs from various generative models.
  • Further investigating interpretable features can help discern the subtle differences between human-written and AI-generated text, especially given the limited availability of such features. Additionally, investigating their vulnerability to adversarial attacks is crucial.
  • Exploring advanced and adaptable learning techniques to effectively address the dynamic and ever-evolving nature of AI-generated text. These methods include adversarial learning, meta-learning, and self-supervised learning (Weber-Wulff et al., 2023).
  • A comprehensive multi-aspect evaluation of detection techniques against adversarial attacks as literature lacks such study to investigate relative vulnerabilities of detection methods. Such an analysis should consider various perspectives, including efficacy across different models and resilience against adversarial attacks. This study can aid in identifying techniques suitable for specific scenarios.
  • Developing hybrid detection strategies that combine features and techniques to enhance robustness and adaptability. For example, one can create a hybrid approach by integrating watermarking and feature-based techniques.
  • A comprehensive understanding of the fundamental feasibility and boundaries as there is often an absence of a thorough grasp of the fundamental feasibility and limitations within current state-of-the-art methods. Therefore, there is a need for deeper exploration and investigation into the theoretical aspects of this task to facilitate the creation of more resilient and efficient techniques while also uncovering novel avenues for research.

### v3

@@ Abstract.
LLM; AI-generated Text Detection; Responsible AI; Watermarking; Paraphrasing Attacks; Data Poisoning
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671463isbn: 979-8-4007-0490-1/24/08ccs: Security and privacy Social network security and privacyccs: Computing methodologies Neural networksccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large Language Models (LLMs) constitute a transformative advancement in the field of natural language processing (NLP). Their applications traverse a wide spectrum of domains, including question answering (Zaib et al., 2021; Bhat et al., 2023), sentiment analysis (Batra et al., 2021; Kheiri and Karimi, 2023) and specially text generation (Senadeera and Ive, 2022). As LLMs are trained on extensive textual corpora, they exhibit a remarkable capacity to produce human-like text.
@@
However, the ubiquity of LLMs brings forth a convergence of opportunities and challenges that necessitate prudent examination. Among these challenges, we encounter the potential for LLMs to produce biased, toxic, or harmful content (Kuchnik et al., 2023; Deshpande et al., 2023; Weidinger et al., 2021; Wen et al., 2023). Additionally, there are concerns related to intellectual property rights infringement (Peng et al., 2023; Stokel-Walker, 2022) and the misuse of LLMs for malicious purposes, such as disseminating misleading information and propaganda (Vykopal et al., 2023; Mozes et al., 2023). These multifaceted considerations highlight the need for judicious evaluation and ethical management of LLMs across different contexts.
Leveraging AI-generated text could be regarded as an efficacious strategy for alleviating the aforementioned challenges. However, the convergence of AI-generated content with human-written texts has reached a point where discerning between the two has become increasingly intricate. The task of distinguishing LLM-generated text from human-written content presents a dual challenge. On one hand, identifying disparities can enhance the quality of AI-generated material. On the other hand, this endeavor also complicates the identification process. In recent years, researchers have proposed various methodologies for detecting AI-generated text to contribute to the ongoing exploration of this landscape.
@@
The rest of this paper is organized as follows: In section 2, we discuss the potential risks arising from the misuse of LLMs. Section 3 provides a comprehensive categorization of various text detection strategies, emphasizing their role as mitigation techniques. Subsequently, in Section 4, we describe the vulnerabilities inherent in these strategies. Moving forward, Section 5 engages in a theoretical exploration of the feasibility of detection, followed by section 6 where we propose new avenues of research. Finally, in section 7, we conclude. Figure 1 demonstrates an overview of the paper.
@@ 2.1. Discrimination, Toxicity, and Harms
A recent study by Wen et al. (Wen et al., 2023) investigates how LLMs generate implicit toxic content that conventional toxicity classifiers struggle to detect. The proposed approach leverages reinforcement learning to uncover and expose this implicit toxicity, emphasizing the need to fine-tune classifiers using annotated examples derived from the attack method for enhanced detection capabilities.
@@ 2.2. Factual Inconsistency and Unreliability of AI Responses
To mitigate such mistakes, various strategies have been proposed through fine-tuning (Lewkowycz et al., 2022; Rajani et al., 2019; Zelikman et al., 2022), prompt engineering techniques like verification, scratchpads  (Cobbe et al., 2021; Nye et al., 2022), Chain-of-Thought (CoT) (Wei et al., 2022), Reversing Chain-of-Thought (RCoT)  (Xue et al., 2023), RLHF (Ziegler et al., 2019; Christiano et al., 2017), iterative self-reflection (Shinn et al., 2023; Madaan et al., 2023), self-consistency (Wang et al., 2023c), society of minds strategy (Du et al., 2023), pruning truthful datasets (Christiano et al., 2023), adjusting the system parameters to limit model creativity (Muneeswaran et al., 2023), external knowledge retrieval (Guu et al., 2020) and training-free methods with likelihood estimation (Kadavath et al., 2022).
The existing body of research emphasize the unspeakable capabilities of LLMs, while also recognizing their proneness to errors. Thus, it is essential to approach LLM outputs with caution.
@@ 4. Vulnerabilities of Detection Techniques
Despite various methods mentioned above, most categories of detection strategies are susceptible to paraphrasing or spoofing attacks. To tackle this, retrieval-based detectors have been employed as detection strategies but these detectors store the output of LLMs in a database and perform semantic searches to extract the best matches. Although it enhances the detector’s resilience against paraphrasing attacks, there are privacy concerns related to storing user-LLM conversations. Additionally, this technique is ineffective when dealing with recursive paraphrasing (Sadasivan et al., 2024).
Moreover, researchers have discovered that by effectively optimizing prompts, LLMs can successfully evade many detection techniques. For example, in a recent work (Lu et al., 2023), Lu et al. propose a novel method called Substitution-based In-Context example Optimization (SICO). In SICO, discriminating features are extracted from both human and AI-generated texts. These features, along with a paraphrasing prompt, are combined and fed to the LLM to modify its output. The prompt is carefully optimized through word and sentence replacements, aiming to minimize the chances of detection while maximizing the similarity between human and AI-generated texts. Evaluation results strongly demonstrate the vulnerability of existing methods.
@@ 5.2. Possibility via Sample Abundance
Interestingly, Chakraborty et al. demonstrate that the AUROC curve, as proposed by Sadasivan et al., might be overly conservative for practical detection scenarios. Specifically, they introduce a hidden possibility by replacing $TV(\mathcal{M},\mathcal{H})$ with $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ in AUROC equation, where $m^{\bigotimes n}:=m\bigotimes m\bigotimes\dots\bigotimes m$ (n times) denotes the product distribution over sample set $\mathcal{S}:=\{s_{i}\},i\in\{1,\dots n\}$, as does $h^{\bigotimes n}$. Thus, for any detector $\mathcal{D}$, with a given collection of i.i.d. samples either from human or AI, the AUROC curve is defined as:
@@
where $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})=1-exp(-nl_{c}(m,h)+o(n)$ and $l_{c}(m,h)$ is the Chernoff information.
Since $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ is an increasing sequence, it eventually converges to $1$ as the number of samples for each distribution increases. Thus, the upper bound of AUROC increases exponentially with respect to the number of samples. It is clear that if this happens, the total variation distance approaches $1$ quickly, and hence increasing the AUROC. Interested readers are encouraged to refer to (Chakraborty et al., 2023) for further details.
@@ 5.3. Impossibility of Strong Watermarking
6. Limitations and Future Research
@@
Curating diverse and representative datasets
is crucial for training and evaluating detection models. The existing datasets may not comprise all the potential types and sources of AI text. Specially, it is essential to curate a dataset containing outputs from various generative models.
Further investigating interpretable features
can help discern the subtle differences between human-written and AI-generated text, especially given the limited availability of such features. Additionally, investigating their vulnerability to adversarial attacks is crucial.
Exploring advanced and adaptable learning techniques
to effectively address the dynamic and ever-evolving nature of AI-generated text. These methods include adversarial learning, meta-learning, and self-supervised learning (Weber-Wulff et al., 2023).
A comprehensive multi-aspect evaluation of detection techniques against adversarial attacks
as literature lacks such study to investigate relative vulnerabilities of detection methods. Such an analysis should consider various perspectives, including efficacy across different models and resilience against adversarial attacks. This study can aid in identifying techniques suitable for specific scenarios.
Developing hybrid detection strategies
that combine features and techniques to enhance robustness and adaptability. For example, one can create a hybrid approach by integrating watermarking and feature-based techniques.
A comprehensive understanding of the fundamental feasibility and boundaries
as there is often an absence of a thorough grasp of the fundamental feasibility and limitations within current state-of-the-art methods. Therefore, there is a need for deeper exploration and investigation into the theoretical aspects of this task to facilitate the creation of more resilient and efficient techniques while also uncovering novel avenues for research.
@@ 7. Conclusions
8. Acknowledgements
This study represents independent research conducted by the authors and does not necessarily represent the views or opinions of any organizations. We express our gratitude to the anonymous reviewers for their valuable and constructive feedback.